OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6, 2) and (7, 4)

OpenStudy (anonymous):

you need to use the slope equation \[(y-y)=m(x-x)\] (2-4)=m(6-7) solve for m

OpenStudy (anonymous):

2

OpenStudy (anonymous):

-2=-1m m=2 yeap